# Convex Approximation of Probabilistic Reachable Sets from Small Samples Using Self-supervised Neural Networks

### v1

@@ V-C Generating speed
In addition to comparing the accuracy, we compare the generating time required to generate useful results. Fig. 8 illustrates the average time cost for generating effective outcomes. The results clearly demonstrate that our method significantly outperforms the baseline approaches, achieving execution times that are an order of magnitude faster. The result supports that the Advantage 2 of the proposed method, NN can generate useful PRS very fast.
@@ VI Conclusion and Future
In this paper, we propose using self-supervised learning to train a transformer-based neural network to generate convex approximations of probabilistic reachable sets. Once trained, the neural network can robustly generate convex approximations that cover the desired percentage of points sampled from the distribution of the probabilistic reachable set only given a few samples from the distribution. The primary advantage of the proposed method is its extremely fast generation time while maintaining performance comparable to that of the baseline methods. Meanwhile, training is also very simple because training inputs can be generated very fast. Only a few experts are required for the labels. No physical information about the system such as disturbance is required.
@@ Acknowledgment
The authors would like to acknowledge the support from the National Science Foundation under Grants CCF-2402689 and CMMI-2138612. Any opinions, findings, conclusions, or recommendations expressed in this paper are those of the authors and do not reflect the views of NSF.

### v2

@@ V Generating speed analysis
Generating time is crucial in deciding whether a method can be used in online applications. Therefore, we compare the time to generate one correct probabilistic obstacle space by each method to prove our method is more useful in online applications. Fig. 9 illustrates the average time cost for generating effective outcomes. The results clearly demonstrate that our method significantly outperforms the baseline approaches, achieving execution times that are an order of magnitude faster. The result supports that the Advantage 4 of the proposed method, NN can generate a useful probabilistic obstacle space very fast.
@@ VI Conclusion and Future
In this paper, we propose using supervised learning to train a transformer-based neural network to generate a confidence region of the real obstacle space distribution, which we define as the probabilistic obstacle space. Once trained, the neural network can robustly generate vertices of the polygon that cover the desired percentage of the real obstacle region, only given a few samples. The primary advantage of the proposed method is its extremely fast generation time while maintaining performance comparable to that of the baseline methods. Meanwhile, training is also very simple because training inputs can be generated very fast. Only a few experts are required for the labels. No dynamic information about the obstacle such as disturbance is required.
@@ Acknowledgment
The authors would like to acknowledge the support from the National Science Foundation under Grants CCF-2402689. Any opinions, findings, conclusions, or recommendations expressed in this paper are those of the authors and do not reflect the views of NSF.